How do you find the slope and y intercept of equations like this! I'm trying to help my 8th grade daughter with algebra and I'm struggling is been way too long since i took algebra. Thanks in advance.
Y=(2-a)x+a

Answer by Theo(13342)

the equation is y = (2-a) * x + a

(2-a) is the slope and a is the y-intercept.

that's because the equation is in the slope intercept form of y = mx + b, where m is the slope and b is the y-intercept.

if you know the value of a, then you can solve the equation for a specific line.

you will get a different equation depending on the value of a.

for example:

if a = 0, then y = (2-a)x + a becomes y = 2x

if a = 5, then y = (2-a)x + a becomes y = (2-5)x + 5 which becomes y = -3x + 5

if a = -5, then y = (2-a)x + a becomes y = 2-(-5))x - 5 which becomes y = 7x - 5

it all depends on the value of a which is not specified, so your equation of y = (2-a)x + a tells you that the slope is (2-a) and the y-intercept is a.

if you know what the answer should be, then hopefully it's what i told you.

if it isn't, then send me the answer and i'll try to figure out what they were really asking.

this equation appears to be fairly simple on the surface.

you just need to know the equation is in slope intercept form and then simply find the slope and the y-intercept.

the general form is y = mx + b
m is the slope
b is the y-intercept
in the equation of y = (2-a)x + a:
(2-a) is the slope.
a is the y-intercept.